What is the concentration of compound Y in an unknown solution if the solution has an absorbance of 0.846 in a glass cuvette with a path length of 0.2 cm? The molar absorptivity of compound Y is 54.2 cm-1(mg/ml)-1 under the conditions used for the absorption measurement.
Answer #1

A = e*l*c

Where

A is absorbance

l is path length

c is concentration

e is molar absorptivity

Given:

A = 0.846

l = 0.2 cm

e = 54.2 cm-1.(mg/mL)-1

Use:

A = e*l*c

0.846= 54.2*0.2*c

c = 7.804*10^-2 mg/mL

Answer: 7.80*10^-2 mg/mL